### v1

# Attention-based Class-Conditioned Alignment for Multi-Source Domain Adaptive Object Detection

Abstract
Domain adaptation methods for object detection (OD) strive to mitigate the impact of distribution shifts by promoting feature alignment across source and target domains. Multi-source domain adaptation (MSDA) allows leveraging multiple annotated source datasets, and unlabeled target data to improve the accuracy and robustness of the detection model. Most state-of-the-art MSDA methods for OD perform feature alignment in a class-agnostic manner. This is challenging since the objects have unique modal information due to variations in object appearance across domains. A recent prototype-based approach proposed a class-wise alignment, yet it suffers from error accumulation due to noisy pseudo-labels which can negatively affect adaptation with imbalanced data. To overcome these limitations, we propose an attention-based class-conditioned alignment scheme for MSDA that aligns instances of each object category across domains. In particular, an attention module coupled with an adversarial domain classifier allows learning domain-invariant and class-specific instance representations. Experimental results on multiple benchmarking MSDA datasets indicate that our method outperforms the state-of-the-art methods and is robust to class imbalance. Our code is available here.
Keywords:
Object Detection Multi-Source Domain Adaptation
1 Introduction
Given the recent advancements in deep learning, deep object detectors have shown impressive results on many benchmark problems ranging from closed-world fixed class settings to open-world scenarios [5, 41, 30, 19]. However, their performance typically degrades when there is a domain shift between the training (source) and test (target) data distributions. To address this challenge, several unsupervised domain adaptation (UDA) methods [16, 25, 4, 10] have been proposed to improve OD performance in the presence of a domain shift.
The conventional UDA methods assume that the source data comes only from a single distribution. In real-world scenarios, the source dataset often comes from multiple distributions with considerable domain shifts due to weather, geographic location, and illumination. This setting is referred to as Multi-Source Domain Adaptation (MSDA) for OD. MSDA is challenging as it requires dealing with discrepancies among sources along with source and target discrepancies. However, having source data from multiple distributions results in enhanced adaptation, increased generalization, and improved robustness against domain shifts[21, 38]. Inspired by the UDA techniques, the typical approach for MSDA involves learning task-specific knowledge from the labeled source datasets, simultaneously trying to make the feature-space domain-invariant that generalizes well to the target domain. Several MSDA methods have been proposed for classification [28, 39, 1], yet MSDA for OD still remains relatively unexplored. This is because in the presence of multiple distribution shifts establishing the trade-off between learning domain-invariant features while preserving task-specific knowledge [15] is challenging when performed both at the image-level and instance level (for OD) compared to only at the image-level (for classification).
(a) No class-wise as in [35, 32]
(b) Prototype based [3]
(c) Ours
Figure 1: Comparision of alignment of different MSDA techniques. (a) DMSN and TRKP implement pairwise-alignment of source-target pair without considering class-wise alignment. (b) Prototype based alignment (PMT) learns specific prototypes to represent each class and domain. Then, the same class prototype for different domains are merged in class-conditioned (domain invariant) prototypes. (c) In contrast, our ACIA learns instance-level domain invariant features by conditioning an attention module to attend a given class.
(a) No class-wise as in [35, 32]
(b) Prototype based [3]
(c) Ours
Figure 2: T-SNE projection of the class distributions on BDD100K of source and target data. (a) Projection without class-specific alignment as in DMSN and TRKP. The classes are not aligned between the two domains. (b) Prototype-based alignment (PMT) uses prototypes for class alignment. Most classes are well aligned but the bike, as it is underrepresented, is not aligned. (c) Our ACIA uses an attention-based adversarial class alignment and manages to align all classes.
Only three methods have been proposed for MSDA of OD - Divide and Merge Spindle Network (DMSN) [35], Target Relevant Knowledge Preservation (TRKP) [32], and Prototype Mean Teacher (PMT) [3]. DMSN and TRKP focus on pairwise alignment between each source and target pairs as shown in Fig. 0(a). This ensures the alignment of each source domain with the target domain. However in a MSDA setting, there are considerable variations in object appearances across domains, so it’s also essential to focus on the class alignment [34]. Else it would lead to class misalignment across domains, which can be observed in Fig. 1(a) [^1] As both methods didn’t provide code or trained model, we used our method without instance-level alignment to generate the TSNE..
Later, PMT proposed a method that uses prototypes for class-conditioned alignment. They use class-specific local and global prototypes to explicitly ensure within-domain and across-domain alignment, as summarized in Fig. 0(b). However, this method has a couple of critical limitations. Firstly, they make a strong preassumption that a single prototype can represent the true distribution of a class. But, generally, a class representation is not isotropic, which can be seen in the Fig. 1(b). Secondly, they rely on noisy pseudo-labelled target domains to learn the prototypes, resulting in class-misalignment due to error accumulation from the pseudo-labels [6] (we showed that in Tab. 5(right)). This issue is particularly problematic for underrepresented classes, which have fewer samples and can lead to misaligned class representations (emperical result in Tab. 6). This misalignment can be observed in Fig. 1(b), where the Bike class (red) is not aligned across the source and target domains. In this scenario, if we calculate the global prototype, it would lie somewhere in the middle of these two distributions, which does not accurately represent the actual class distribution. Consequently, this results in the degradation of the detector’s performance.
This highlights the need for a better class-conditioned instance alignment mechanism in MSDA for OD. In this paper, we propose an Attention-based Class-wise Instance Aligner (ACIA) for MSDA, aiming at efficiently aligning class instances across domains, without relying on prototypical representation. The alignment process of ACIA is illustrated in Fig. 0(c) where each object category is aligned together across domains. In ACIA, the instances of each ROI-pooled features are infused with the class information by an attention operation. The class information is learned by using a label embedding. To compute the attention output, the query and key values are obtained from ROI-Pooled features, and the key is obtained from label embeddings. The output of the attention is used by adversarial domain discriminatior. The gradients are reversed from the output of the instance-level feature extractor forcing those features to be domain-invariant and class-wise aligned at the same time. Our method relies on just a single class embedding to store relative class information for each object category across domains. This allows a simple yet efficient multi-source class-conditional alignment. The benefits of our class-conditioned aligner can also be observed from Fig. 2, where our ACIA provides better class alignment across domains and robustness to imbalanced data.
Our main contributions are summarized as follows.
(1) An MSDA method called ACIA is proposed that integrates a conceptually simple class-wise instance alignment scheme. Instance alignment is implemented with an attention mechanism where instances from all domains are aligned to the corresponding class by the attention component.
(2) The combination of attention and adversarial domain classifier provides instance representations that are simultaneously domain-invariant and class-specific.
(3) Extensive experiments are conducted on three benchmark MSDA settings to validate the proposed ACIA alignment and ablation of the components. ACIA achieves state-of-the-art performance on all three benchmarks while remaining robust when adapting to imbalanced data. It is the simplest class-conditioned alignment process among the MSDA methods for OD.
2 Literature Review
1. Unsupervised Domain Adaptation. UDA methods aim to learn a model that performs well on the target domain by combining the information from the labeled source data and unlabeled target data. A common approach for UDA is using an adversarial discriminator to generate domain-invariant representation. [8] integrated a GRL-based (reverses the gradient during backpropagation) discriminator at the image and instance level to align the features across domains. Later, [24] showed that strong and weak alignment of high-level and low-level features can improve the performance in UDA. [42] emphasizes the fact that in OD not all parts of the image are equally important. So, they focus on aligning the parts of the images that are more important using clustering and adversarial learning. [40] argues that aligning foreground regions is more important than background regions. They used the output of RPN as an attention map and prototypes for instance-level alignment. [14, 7, 23] tried to mitigate domain shift at the image level. They used Cycle-GAN to convert the source domain to the target domain and minimize the domain shift. These works focus on a single source, however, in this paper, we focus on the MSDA setting where the source data itself has multiple distribution shifts.
2. Multi-Source Domain Adaptation. MSDA methods seek to adapt a model by using samples from multiple labeled source datasets and unlabeled target data. Only three works have been proposed to solve the problem of MSDA in OD - DMSN [35], TRKP[32], and PMT[3]. [35] learns pseudo-subnets for each domain and uses adversarial alignment to align each source subnet with the target. The weight of the target subnet is obtained as the weighted exponential average of the student subnets. [32] proposed an architecture with one detection head for each source domain. These detection heads are trained by using adversarial disentanglement for cross-domain alignment. They also introduced a mining strategy to weigh each source image based on its relevance to the target domain. Later [3] propose to use class and domain-specific prototypes as a representative of each class per domain. These prototypes are aligned by using contrastive loss to achieve class-conditioned adaptation. Out of the three methods only [3] focuses on class-wise alignment. However, they rely on the assumption that class representations are isotropic and also they suffer from pseudo-label error accumulation[6]. In contrast, our ACIA uses an attention-based adversarial aligner that is immune to pseudo-label error accumulation.
3. Instance-Level Alignment in Domain Adaptation. The instance-level alignment is applied in a DA setting to ensure domain-invariant features at the instance level. [8] uses an adversarial domain classifier at the instance level similar to the image-level discriminator. [13, 33] improved on the work of [8] by giving weight to each ROI-pooled feature during instance-level alignment. This ensures that more importance is given to hard-to-classify features. [42] uses a COGAN to regenerate the ROI-pooled features and later align them using adversarial training. These works align the instance-level features in a class-agnostic way. [29] emphasizes the importance of object category-wise alignment. For this, they used several memory bank prototypes to store class information across domains and they are aligned by using class-specific discriminators. [37] learns a prototype for foreground and background across source and target domain. These prototypes are aligned to produce better predictions on the target domain. [34] learns graph-induced prototypes and proposes a loss to align them. [40] used prototypes and proposed to do instance-level alignment in a class-wise fashion using a prototype-based method. Nevertheless, all these methods rely on learning prototypes for instance-level class alignment in OD, based on the assumption that class representation is isotropic and used pseudo-labels for alignment.
3 Proposed Method
The MSDA setting consists of $N$ source domains $S_{1},S_{2},…,S_{N}$ and the target domain $T$. Similar to UDA setting, the object categories across all domains are fixed to a pre-defined set of categories and the number of object categories is $K$. Each source domain $S_{j}$ consists of images and corresponding labels, denoted as $S_{j}=\{(x^{j}_{i},y^{j}_{i})\}^{M_{j}}_{i=1}$, where $j=1,2,...,N$, and $M_{j}$ is the number of images in the source domain $j$. Here, $x^{j}_{i}$ is the $i$th image from source domain $S_{j}$ and $y^{j}_{i}$ is the set of object labels and corresponding bounding boxes. The target domain $T$ only has images, with no bounding boxes and class labels information. It can be represented as $T={\{x_{i}^{T}\}^{M_{T}}_{i=1}}$, where $M_{T}$ is the number of images in the target domain and $x_{i}^{T}$ is the $i$th image from the target domain.
The training framework of our proposed ACIA is shown in Fig. 3. Our model is based on the Mean-Teacher (MT) framework[10]. The components responsible for image-level and instance-level alignment are integrated into the MT framework. They are the multi-source discriminator and attention-based class-conditioned instance-level discriminator. The MT training process is performed in two stages. First, the student model is trained in a supervised manner on all source domain data, typically called the burn-in stage in the MT training process [10]. After this step, the teacher model weights are initialized as an exact copy of the student model weights. Then, mutual learning of teacher-student begins on both the labeled source datasets and unlabeled target data. In the case of source datasets, the same supervised training is continued. For the target dataset images, two augmented versions are created for each image, called the weak and strong augmentation. The teacher model receives the weak augmented version and it computes the pseudo-labels. The student model receives the strong augmented version and computes its loss with the pseudo labels provided by the teacher as ground truth. Along with that, the multi-source discriminator and class-conditioned instance-level discriminator are trained for domain-invariant feature generation and class-wise alignment respectively. The weight of the teacher model is the exponential moving Average(EMA) of the weight of the student model. At the time of inference, only the teacher model is used. The rest of the section describes all the loss functions.
Figure 3: Overview of the training architecture of our ACIA. The overall architecture is a teacher-student model in which the student learns from multiple sources of data. An image-level classifier is introduced to globally align features, and an instance classifier to align instances in a class-conditional way, through a transformer-based attention block (see white boxes).
1. Source Data Training The source data from all the domains have annotations available. Thus the standard supervised training is applied for the source datasets where the student network weights are updated by backpropagating the detection loss:
$\mathcal{L}_{\mbox{sup}}=\sum_{j=1}^{N}\sum_{i=1}^{M_{j}}\mathcal{L}_{\mbox{cls}}(x^{j}_{i},y^{j}_{i})+\mathcal{L}_{\mbox{reg}}(x^{j}_{i},y^{j}_{i})\ ,$ (1)
where $\mathcal{L}_{\mbox{reg}}$ is the smooth-L1 loss [22] used for the bounding box regression and $\mathcal{L}_{\mbox{cls}}$ is the focal loss [17] used for bounding box classification. The source datasets are used in both the training stages, ie, the burn-in and teacher-student mutual learning.
2. Target Data Training Target datasets are used in the training process at the teacher-student mutual learning stage. For every image from the target domain, two augmented versions are generated - strong and weak augmentation. For the strong augmentation, we used color jittering, grayscale, Gaussian blur, and cutout patches. For weak augmentation, image rescaling and horizontal flips are used. The weak augmented image is fed to the teacher model and the output is post-processed to generate pseudo-labels $\widetilde{y}$. The post-processing applied is the confidence thresholding [10]. Let the output of the teacher network for an image $x_{i}$ be $\hat{y}_{i}=\{c_{i}^{j},p_{i}^{j},b_{i}^{j}\}_{j=1}^{D}$ where $c_{i}^{j},p_{i}^{j},b_{i}^{j}$ are predicted class, probability and box respectively and $D$ is the number of output detections. Then $\widetilde{y}_{i}=\{c_{i}^{j},b_{i}^{j}|p_{i}^{j}>\theta\}$ where $\theta$ is a threshold used for filtering noisy pseudo-labels. The student network then receives the strongly augmented version and computes its loss. The loss function used is the same as eqn. 1, except the pseudo-labels are employed:
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x_{i},\widetilde{y}_{i})+\mathcal{L}_{\mbox{reg}}(x_{i},\widetilde{y}_{i})$ (2)
Once the student weights are updated via backpropagation, the teacher weights are updated as $w_{t}=\alpha w_{t}+(1-\alpha)w_{s}$ where $w_{t},w_{s}$ are the weights of teacher and student networks respectively and $\alpha$ is a smoothing factor.
3. Image-Level Alignment In the domain adaptation problem, the task-specific knowledge is learned from the source domain. But, for the model to perform well on the target domain, we have to learn a representation that is also domain-invariant. In a UDA setting, this is achieved by attaching a domain discriminator (binary classifier) to the output of the backbone using a gradient-reversal layer (GRL) [16]. This results in a min-max optimization game between the backbone and the domain-discriminator, leading to domain-invariant feature representation. Unlike UDA, in MSDA setting the source data comes from various distributions, so we used a multi-class discriminator instead of a binary discriminator. In our work, we used a CNN-based classifier that outputs a domain prediction map from the last feature map of the backbone. The cross-entropy loss is then applied to all pixels of the prediction map and the results are pooled using global average pooling. The training loss of the image-level discriminator can be represented as:
$\mathcal{L}_{\mbox{dis}}=-\frac{1}{WH}\sum_{j=1}^{N+1}\sum_{i=1}^{M_{j}}\sum_{h=1,w=1}^{H,W}d_{i}^{j}\log(D(F(x_{i}^{j})_{wh})$ (3)
Here, W and H are the width and height of the prediction map. $D(F(x_{i}^{j})_{wh}$ is the output of the multi-class discriminator for pixel $w,h$ of the prediction map. D and F represent the discriminator and Feature extractor (the backbone network in this case) respectively. $d_{i}\in\{0,1\}^{N+1}$ whose value is one for the correct domain of the $i$th image and zero everywhere else.
4. Instance-level Alignment The instance-level alignment is performed in a class-wise fashion in our method. Class-wise alignment helps to align features of objects from the same class regardless of the difference in their visual attributes (such as appearance, scale, orientation, and other characteristics of objects) both within and across domains. Thus even in the presence of class imbalance, underrepresented classes will be focused on ensuring the alignment within and across domains. As shown in fig 2, without class-wise alignment, the separation of class features of the same category across domains is high. PMT [3] is the first MSDA approach in OD that implements class-wise alignment using global and local prototypes. The local prototypes ensured within domain alignment whereas the global prototypes ensured across domain alignment. In contrast to using prototypes for each object category and each domain, we use an efficient attention-based module to ensure class-wise instance alignment.
The instance-level alignment in OD is performed using the region features from each instance. Different from image-level alignment, we have features for a single object category at the instance level. Following the previous approaches, an adversarial discriminator on RoI-Pooled features can be used to perform the instance-level feature alignment. But this alignment process is class-agnostic. If we can perform this adversarial learning in a class-wise fashion, the resulting feature alignment will be class-conditioned. The class information is available for the source domain from the annotations. For the target domain, we can use the filtered pseudo-labels generated by the teacher model. However, the pseudo-labels are noisy, even after filtering. Using them would lead to accumulation error during training from the noises in pseudo-labels [6]. We empirically also found that using the pseudo-labels in class-conditioned alignment harms the model’s performance (see Tab. 5).
To use class information during the alignment process, we create learnable embeddings for each class label. These embeddings can be coupled with region embeddings from the RoI Pool using simple operations like concatenation or multiplication. It is observed that the attention mechanisms are effective at finding relations between two vectors by dynamically focusing on different parts of the sequences [2, 20]. So, we decided to employ an attention-based alignment operation (the results from empirical comparison are available in table 5). For this, two linear projections of the ROI-Pooled features and a linear projection from the label embedding are created. The projections from region embedding are used as query($Q$) and value($V$) respectively, while the projection from label embedding is used as key($K$) for the attention block (see Fig. 3). The attention block uses a multi-head attention as proposed in the original paper [27] which can be expressed as:
$\mbox{Attention}(Q,K,V)=\mbox{softmax}\left(\frac{QK^{T}}{\sqrt{d_{K}}}\right)V\,$ (4)
where $d_{K}$ is the size of the key $K$. The dot-product of $K$ and $Q$ allows our model to learn the relation between the object features and class labels based on object category and image domain. Later, the dot-product of this with $V$, allows our model to learn which part of the object feature to focus on for class-conditioned adaptation. We also combine multiple single-head attention to get multi-head attention. This allows our method to learn multiple such relations. At the end of the attention block, a classification head is used to predict the domain of the input instance. The instance aligner is trained with the cross-entropy loss:
$\mathcal{L}_{\mbox{cdis}}=-\sum_{j=1}^{N}\sum_{i=1}^{M_{i}}\frac{1}{Z_{i}^{j}}\sum_{r=1}^{Z_{i}^{j}}d_{i}^{j}\log(D(R(F(z_{i,r}^{j}))))\ ,$ (5)
where $D$, $R$, and $F$ represent the Discriminator, ROI-Pooling, and Feature extractor of the student model; $Z_{i}^{j}$ is the number of Ground Truth(GT) instances in $i$th image from domain $j$, and $z_{i,r}^{j}$ is the $r$th GT box on $i$th image from domain $j$. The target of the discriminator, $d_{i}\in\{0,1\}^{N}$, is the domain label for each GT box as in Eq. 3. Note that $d_{i}$ is of length $N$ because we are not considering (pseudo) GTs from the target image.
Final Training Objective: After starting the student-teacher mutual learning all the losses discussed above are combined and used for training the student model. So, the final training loss for the student model would be:
$\mathcal{L}=\mathcal{L}_{\mbox{sup}}+\alpha\mathcal{L}_{\mbox{unsup}}+\beta\mathcal{L}_{\mbox{dis}}+\gamma\mathcal{L}_{\mbox{cdis}}\ ,$ (6)
where $\alpha$, $\beta$, and $\gamma$ are hyper-parameters for weighing each losses. The teacher model is updated as an EMA of the student model at each iteration.
4 Results and Discussion
4.1 Experimental Methodology
1. Implementation Details - In our experiments, we follow the same settings as [35, 32, 3]. In all experiments, Faster-RCNN[22] with Imagenet pretrained VGG-16 backbone is used as the detector. Images are resized such that the shorter side of the image is 600 while maintaining the image ratios, following the ROI-alignment[12] based implementation of Faster-RCNN. For the hyperparameters, we set $\alpha$, $\beta$ and $\gamma$ to 1, 0.1 and 0.3, respectively. The pseudo-label threshold was set to 0.7 for all the experiments. As mentioned in 3, our training is two-staged. We trained our model for 15 epochs in both the settings. We set the weight smoothing coefficient for EMA of teacher mode to 0.9996. All experiments were done on 4 A100 GPUs, with batch size and learning rate of 8 and 0.2 respectively. We implemented our method using Pytorch.
2. Baseline Methods - We compared our results against four settings: 1 - Lower Bound - Supervised training on the merged source data. 2 - UDA - Source domains are blended and the UDA method is used. 3 - MSDA - SOTA MSDA methods fall into this category. 4 - Upper Bound - We have 3 upper bounds: Target Only - Supervised training on target data, All-Combined - Supervise training on target and source combined, Fine-Tuning - Trained on source domain, then fine-tuned on the target domain.
3. Datasets and Adaptation Setting For our experiments we use 5 different datasets with various objects and different domains and capturing conditions: BDD100K [36], Cityscapes [9], KITTI [11], MS COCO [18] and Synscapes [31]. The MSDA problem settings used in the empirical study are summarized in Tab. 1. For each setting, the source and target domains used are listed as well.
Setting Src.1 # Img. Src.2 # Img. Src.3 # Img. Target # Img.
Cross-time Day 36,728 Night 27,971 - - Dawn 5,027
Cross-camera Cityscapes 2,831 Kitty 6,684 - - Day 36,728
Mixed Domain Cityscapes 2,975 COCO 71,745 Synscapes 25,000 Day 36,728
Table 1: Summary of the different MSDA settings used in our study.
4.2 Results
1. Cross-time Adaptation - Daytime and Night subsets of BDD100K dataset are used as source domain, while the Dusk/Dawn subset of BDD100K is used as target domain. In this setting the images are captured at different times. This causes a domain shift due to variation in illumination.
We reported the result of our method, against other baseline in Tab. 2. It can be observed that the DAOD methods outperforms the Lower Bound. But, there is a slight decrease in performance when the Night domain is added to the Daytime domain. This is because the UDA Blending methods did not consider the domain shift among the source domains, resulting in a degradation of performance. This emphasizes the importance of research on the MSDA methods. Our method beats the previous state of art MSDA method[3] by 2.6 mAP. Also, our method outperforms Target-Only and All-Combined Upper Bound Settings. The poor performance of the Target Only setting is due to fewer samples in the target domain compared to source domains.
2. Cross-camera Adaptation- Cityscapes and Kitty datasets are used as source domains, while the Daytime subset of BDD100K is used as the target domain. In this setting the images are captured using different cameras. This causes a domain shift due to variations in the camera’s resolution and viewpoint. In this setting, the training and evaluation is done only on the car object category (the only overlapping class across the domains).
The AP on car compared against the baselines is reported in Tab. 2. The general variation of performance in this setting is similar to Sec. 4.2 setting. Our proposed method is able to outperform the previous method by 0.4%. Our result is also very close to the target-only upper-bound. In this setting there is only one object category, thus the class-conditioning has no impact in instance alignment. As shown in the ablation studies, the class-conditioning during instance alignment is crucial for the performance improvement.
Setting Method D N D+N C K C+K
Lower Bound Source Only 30.4 25.0 28.9 44.6 28.6 43.2
UDA Strong-Weak[26] 31.4 26.9 29.9 45.5 29.6 41.9
Graph Prototype[4] 31.8 27.6 30.6 - - -
Cat. Regularization[33] 31.2 28.4 30.2 46.5 30.8 43.6
UMT[10] 33.8 21.6 33.5 47.5 35.4 47.0
Adaptive Teacher[16] 34.9 27.8 34.6 49.8 40.1 48.4
MSDA MDAN[38] - - 27.6 - - 43.2
[…]
DMSN[35] - - 35.0 - - 49.2
TRKP[32] - - 39.8 - - 58.4
PMT[3] - - 45.3 - - 58.7
ACIA (Ours) - - 47.9 - - 59.1
Upper Bound Target-Only - - 26.6 - - 60.2
All-Combined - - 45.6 - - 69.7
Fine-Tuning - - 50.9 - - 72.1
Table 2: Detection mAP of our method compared against the baselines in the cross-time (left) and cross-camera (right) setting. Source domains are daytime (D) and night (N) subsets and the target is Dusk/Dawn subset of BDD100K in cross-time setting. And, for the cross-camera setting source domains are Cityscapes (C) and Kitty (K) and the target is Daytime subset of BDD100K.
3. Extension to Mixed Domain Adaptation - MS COCO, Cityscapes, and Synscapes datasets are used as source domain. While, the Daytime subset of BDD100K is used as target domain. In the real-world data the domain shift is not limited to a single factor. So, in this setting domain shift occurs due to multiple and variable reasons. Also, in this setting, we have three source domains, instead of two. These reasons make this setting challenging and very close to practical setting.
The empirical results from this setting is presented in Tab. 3. In this setting the Lower Bound is performing better than the UDA Blending baseline. We hypothesize this is due to large domain shift among domains so blending all the source domain as one and using UDA is resulting in performance degradation. This highlights the difficulty of this setting. When compared with the previous MSDA methods, our method is outperforming them. When Cityscapes and MS COCO are used as source domain, we’re outperforming the previous SOTA result by 2.3 mAP. When Synscapes is also added to the source domain this number rises to 2.6 mAP. This shows that our method continues to improve the performing when source domain are added. Our method is also outperforming the Target Only Upper Bound.
Setting Method C C+M C+M+S
Lower Bound Source Only 23.4 29.7 30.9
UDA UMT[10] - 18.5 25.1
Adaptive Teach.[16] - 22.9 29.6
MSDA TRKP[32] - 35.3 37.1
PMT[3] - 38.7 39.7
ACIA (Ours) - 41.0 42.3
Upper Bound Target-Only - - 38.6
All-Combined - 47.1 48.2
Fine-Tuning - 49.2 52.5
Table 3: mAP on 7 object categories of our method compared against baselines in the mixed setting. Source domains are Cityscapes(C), MS COCO(M), and Synscapes(S) datasets while the Daytime domain of BDD100K is target domain.
4.3 Ablation Studies
Unless mentioned, all the ablation studies are performed on the cross-time adaptation setting as in [35, 3]. The impact of the components performing image-level and instance-level feature alignment is analyzed along with other possible design choices for class-conditioned instance alignment.
1. Impact of the adaptation components - First, we analyzed the impact of the individual components of our proposed method. These components include image-level discriminator and instance-level discriminator. To better understand the importance of class conditioning in the instance alignment, we further examined the instance classifier with and without class-wise attention. The results are reported in Tab. 4. It can be observed that just by adding the image-level discriminator, the performance improved by a huge margin. Furthermore, by adding the instance-level discriminator (without class-embedding) the performance of the model increased by 3.4% showing the importance of instance-level alignment. The performance is further boosted by integrating the class-information into the instance-level alignment.
Image-LA Instance-LA Instance-LA $\textrm{AP}_{50}$
w/o class-embed. w/ class-embed.
31.7
$\checkmark$ 41.9
$\checkmark$ $\checkmark$ 45.3
$\checkmark$ $\checkmark$ 47.9
Table 4: Impact of individual components our ACIA method. Image-LA: Image-Level Alignment, Instance-LA: Instance-Level Alignment.
2. Design choices for class-wise instance alignment - The instance-level alignment component can be designed in different ways. Apart from the attention-based method, class information can be incorporated to instances by concatenation and multiplication. Tab. 5(left) reports the variation in performance with different design choice. It is clear that attention is the superior design choice. We believe this is because we’re learning just one class-embedding layer across every domain. When we’re merging the class-information to the ROI-Pool features, attention can learn what information is more important for a particular domain in a better way than multiplication and concatenation. This is because of the design of the attention layer, allowing it to learn complex relations [27].
Merging D+N
Concatenation 46.3
Multiplication 46.8
Attention 47.9
Setting With Without
Target Target
Cross-Time 43.3 47.9
Mixed 37.9 41.0
Table 5: (left) Different design choices for integrating class information in instance alignment; (right) Detection performance degrades when target domain is used in the instance alignment process.
3. Impact of target domain on class-wise alignment - Instance-level alignment differs from image-level alignment in that it focuses solely on the foreground region of the image. In this section, we investigate the impact of employing or omitting the target image in instance-level alignment. As reported in Tab. 5 (right), utilizing target data in both scenarios leads to a decline in detector performance. In a DA scenario, the target data lack labels, making it challenging to distinguish between foreground regions and object categories. One approach to address this challenge is by employing pseudo-labels; however, they often introduce noise and degrade performance. Nonetheless, in MSDA, the availability of multiple source domains enables us to choose which domains to align at the instance level. Hence, our work focuses on aligning solely the source domains.
4. Detection Visualization In Fig. 4 we present the visualization of detections on BDD100k cross-time for the three multi-source adaptation approaches presented in the paper: no class-wise adaptation, prototype-based class-conditional adaptation and our attention-based class-conditional adaptation.
5. Level of Imbalance In Tab. 6 we compare our method against PMT for different levels of imbalance. As expected, when the dataset is balanced the two methods perform similarly. However, when the level of imbalance increases, our method obtains much better mAP. This shows that the error accumulation from the pseudo-labels is more problematic in the case of imbalanced datasets. Our class-conditioned attention model solves that issue, as it uses the same model for all classes and therefore can better deal with high levels of imbalance.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) No class-wise (b) Prototype based (c) Ours
Figure 4: Examples of ODs on BDD100k cross-time setting, with different types of instance-level adaptation. (a) ODs with a multi-source adaptation without instance level class-conditional adaptation as in [35, 32]. (b) ODs with the prototype-based class-conditional adaptation [3]. (c) ODs with our ACAI approach.
Method Bike Bus Car Motor Person Rider Light Sign Truck mAP
Mostly Balanced 25.0 50.0 10.0 15.0
PMT 43.8 52.3 - 28.0 - 34.4 - - - 39.6
Ours 45.8 48.9 - 30.7 - 31.0 - - - 39.1
Slightly Imbalanced 4.6 9.2 1.8 59.6 2.8 22.0
PMT 42.1 52.7 - 25.1 43.1 34.0 - - 50.5 41.2
Ours 44.3 54.2 - 25.3 45.3 40.6 - - 50.4 43.4
Mostly Imbalanced 1.2 2.4 0.5 15.4 0.7 32.8 42.3 5.7
PMT 49.2 56.2 - 29.6 49.9 32.0 43.2 55.8 55.3 46.4
Ours 52.1 59.2 - 29.2 53.1 36.1 46.2 58.1 57.1 48.9
Highly Imbalanced 0.5 1 57.8 0.2 6.5 0.3 13.4 17.8 2.4
PMT 55.3 59.8 67.6 29.9 47.6 32.7 46.3 56.0 57.7 50.3
Ours 56.1 61.0 69.2 31.9 51.8 39.8 49.2 59.0 61.0 53.2
Table 6: Detection results per-class on cross-time setting for PMT and our method with different levels of imbalance. The dataset is quite imbalanced. To create a more balanced dataset we reduced the number of classes by removing those that are creating the imbalance. For each level of imbalance, we report the relative importance of each class in percentage.
5 Conclusion
Recent studies in UDA have observed that doing instance alignment in a class-conditioned fashion has impressive results. However, a comprehensive study on class-conditioning in MSDA for OD to understand its impact and contribution for instance alignment is missing in existing works. Also, the existing techniques for class-conditioning have several limitations that can lead to misalignment of instances. We empirically observed this misalignment in the case of underrepresented classes. Our work is addressing this research gap with the goal of developing an efficient class-conditioned instance alignment for MSDA. We propose an attention-based class-conditioning scheme for instance alignment where attention operation between region embeddings and label embeddings coupled with an adversarial domain classifier promotes the domain-invariance of object instances in a class-conditioned manner. The resulting instance-alignment scheme obtained state-of-the-art results on popular MSDA benchmarks while being robust to class imbalance.